# TactileNet: Bridging the Accessibility Gap with AI-Generated Tactile Graphics for Individuals with Vision Impairment

### v1

@@ 4 Experimental Settings
4.1 Fine-tuning Configuration
We fine-tune the Anything V3 AnythingV3 (2024) across 66 distinct classes. Each class, supported by training images ranging from a minimum of 9 (llama class) to a maximum of 102 (helicopter class), with an average of approximately 16 images per class. We levearge the sourced tactile images along with the prompts to enable the development of class-specific adapters. These lightweight modules specialize in generating tactile graphics for each category and are integrated atop frozen versions of SD v1.5 for image generation.
@@ 4.1 Fine-tuning Configuration
For fine-tuning, the DreamBooth configuration prioritized prior preservation with a loss weight of 1.0, ensuring the model retains its ability to generate generic images while adapting to tactile-specific features. The LoRA setup used a network module with linear dimensions set to 32 and an alpha parameter of 16, which controls the scaling of low-rank updates. This configuration ensures effective parameter adaptation without requiring pre-loaded weights.
4.1.1 Optimization and Hardware Configuration Parameters
@@
The optimization process employed the AdamW8bit optimizer Kingma and Ba (2017), renowned for its efficiency, particularly with a learning rate set at $1\times 10^{-4}$ for the UNet Ronneberger et al. (2015) and $5\times 10^{-5}$ for the text encoder, operating under a constant rate scheduler without warm-up steps. Our experiments for fine-tuning were facilitated by the NVIDIA Tesla T4 GPUs provided by Google Colab. A training batch size of 6 was used, utilizing mixed precision training at FP16 across a maximum of 20 training epochs.
4.2 Image Generation Configuration
@@
The primary configuration for text-to-image generation involved the DPM++ 2M Karras sampling method, with 20 steps at image dimensions of 512×512 pixels at a CFG scale of 7. For image-to-image translation experiments, configurations were similar, except for an additional denoising strength of 0.9 to better retain the structural essence of the original images while incorporating new tactile features.
@@
For some challenging classes, deviations were necessary to ensure optimal structure and clarity. Objects requiring finer details, such as bee wings, elephant textures, and duck feathers on water, utilized higher denoising strengths (0.96–1.0) and adjusted CFG scales (8–10) to improve contrast and edge clarity. Classes with intricate internal features like basketball, dinosaur T-Rex teeth, and camera lenses benefited from additional lineart-based ControlNet modules to refine their contours. Water-based objects, such as the sailboat and floating duck, incorporated negative prompts to suppress excessive water details, ensuring the main subject remained distinct. For simpler structures like egg, book, and hatchback car, a LoRA-only approach sufficed, requiring no ControlNet assistance. We utilize NVIDIA TITAN V GPUs with 12288 MiB of memory for generating the images.
4.3 Evaluation Protocol
@@
To assess the quality and usability of the generated tactile graphics, we developed a custom evaluation protocol in collaboration with tactile graphics designers. Standard metrics such as Fréchet Inception Distance (FID) Heusel et al. (2017) and Structural Similarity Index (SSIM) Wang et al. (2004) are insufficient for this specialized domain, as they do not account for the unique requirements of tactile graphics, such as line clarity, texture, and adherence to accessibility guidelines BANA (2025). Instead, our protocol prioritizes expert evaluation by our industry partners, ensuring that the generated graphics meet the practical needs of visually impaired users.
4.3.1 Interface Design and Rationale
Our evaluation interface presents evaluators pairs of images:
@@ 4.3.1 Interface Design and Rationale
  • Tactile Image: The generated tactile graphic or a sourced tactile graphic from established libraries (e.g., APH, Perkins).
This side-by-side comparison allows evaluators to assess how well the tactile graphic captures the essential features and posture of the subject while adhering to tactile design principles. Evaluators are blind to the source of the tactile graphic (generated vs. sourced), ensuring an unbiased assessment.
4.3.2 Evaluation Questions and Metrics
@@
Assess whether the tactile graphic follows established tactile graphics standards (e.g., BANA guidelines BANA (2025)).
@@
4.3.3 Natural Image Selection and Prompt Generation
@@
  • Original Prompt: The prompt generated during the prompt generation phase (Section 3.2.2).
  • Paraphrased Prompt: A refined version of the original prompt, paraphrased using DeepSeek DeepSeek (2025) with the instruction: ”Paraphrase the given prompt for tactile generation while preserving the ’tactile’ subject and essential features.”
@@
4.3.4 Statistical Analysis Plan
@@
5 Results and Discussion
In this section, we present results including TactileNet dataset statistics (Table 2) and outcomes from text-to-image and image-to-image translation tasks.
5.1 Image-to-Image Translation Evaluation
Table 1 summarizes the quality ratings for both sourced and generated tactile images. Figure 1 illustrates some samples from our evaluation: the first row presents the natural reference images, the second row displays the sourced tactile graphics, and the third row showcases the tactile graphics generated using our fine-tuned adapters.
Table 1: Quality Ratings for Generated vs. Sourced Tactile Graphics
Category Generated (%) Sourced (%)
Accept as Is 32.14 35.71
Accept with Minor Edits 39.29 39.29
Accept with Major Edits 28.57 21.43
Reject (Useless) 00.00 3.57
5.1.1 Key Findings
Our evaluation reveals that both sourced and generated tactile graphics achieved 100% alignment with natural images in terms of pose and structural features (Q1), confirming their accuracy in representation. However, adherence to tactile accessibility guidelines (Q2) was slightly higher for sourced graphics (96.43%) compared to generated ones (92.86%), suggesting that while our fine-tuned models perform well, further refinements are needed for optimal accessibility.
Additionally, 28.57% of generated graphics required major edits, primarily due to excessive visual complexity, such as unintended 3D perspectives in objects like Chair and Bed (Figure 1), which can hinder tactile interpretability. A small fraction of sourced tactile images (3.57%) were rejected mostly due to human-induced mismatches in dataset curation, where incorrect natural images were paired with tactile graphics, as seen in the camel species (Figure 4). This analysis highlights the strengths of AI-generated tactile graphics while emphasizing key areas for improvement, particularly in simplifying complex structures and refining dataset sourcing for higher semantic accuracy.
@@ 5.1.1 Key Findings
Table 2: Overview of the dataset statistics: scrapped tactile images (Source) used for training the models and the output tactile images (Generated) of our text-to-image adapters.
5.2 Text-to-Image Translation Results
We generated a total of 32,000 tactile images across 66 classes, with 7050 images retained after an initial non-expert human filtering. Unlike the image-to-image translation setup, no natural image was provided as input; instead, the model relied solely on textual prompts to generate tactile graphics.
Figure 5: Prompt editing for customization: (Left) T-shirt generated using the base prompt; (Center) Logo removed by adding ”logo” as a negative prompt; (Right) Pocket removed by omitting the keyword ”pocket” from the prompt.
@@ 5.2 Text-to-Image Translation Results
The per-class count is summarized in Table 2. Our fine-tuned adapters demonstrated adaptability through prompt-based edits (Figure 5), enabling modifications such as logo removal or pocket additions. The prompt used for generating images was: Base Prompt: Create a tactile graphic of a t-shirt with a pocket for the visually impaired, highlighting the round neckline, pocket edges, and hem with raised lines. Ensure the pocket is a distinct raised rectangle, allowing users to discover and feel the detail on the shirt’s chest.
6 Conclusion
@@
This study marks a significant advancement in the development of AI-driven accessibility tools through the application of Generative Artificial Intelligence, effectively addressing the critical need for scalable, high-quality tactile graphics. Our contributions are threefold:
@@
  • TactileNet Dataset: A first-of-its-kind dataset of 1,029 expert-curated tactile images across 66 classes, enabling AI model training tailored to tactile design principles.
@@
  • Efficient Fine-Tuning: Integration of LoRA and DreamBooth reduces labor costs while achieving 92.86% adherence to tactile guidelines, rivalling manually sourced graphics.
@@
  • Human-Centric Evaluation: A protocol validated by tactile experts ensures usability, with generated graphics requiring no rejections and matching natural images in alignment 100%.
@@
By automating tactile graphic generation, our method accelerates production and democratizes access to educational materials for visually impaired learners. This work illustrates how AI can drive social good, providing a blueprint for inclusive technology that prioritizes human needs. We encourage further research and collaboration in this area to enhance accessibility and empower visually impaired individuals.

### v2

@@ IV EXPERIMENTAL SETTINGS
IV-A Baseline: Instruction-Tuned Stable Diffusion (InstructPix2Pix)
As a baseline, we fine-tuned Stable Diffusion (SD) [7] using the InstructPix2Pix method [30, 31] on 1,029 triplets consisting of natural images, sourced tactile graphics, and corresponding edit prompts. The natural images were gathered by a team of five undergraduate students, who reviewed each sourced tactile graphic and selected the best-matching natural image from online sources.
The edit prompts were deliberately kept simple and generic—e.g., “Create a tactile image of the given natural image of a [class].”—to align with the original instruction-tuned training paradigm. Over 100 such generic prompts were created using ChatGPT and randomly paired with images, with [class] replaced by the object name. We also experimented with complex, template-based prompts detailing features and patterns. In both cases, however, the model exhibited significant overfitting: during testing, it disregarded input images and memorized the prompt-class pairs from the training data. Changing prompts often resulted in hallucinated or irrelevant outputs, highlighting the baseline’s strong dependence on text rather than visual conditioning.
To increase data diversity and mitigate overfitting, we expanded the dataset by pairing each of the 1,029 tactile-natural image pairs with the full set of 100 generic prompts, resulting in 102,900 training examples. Nevertheless, the baseline underperformed on unseen test images, defaulting to memorized prompt outputs rather than adapting to new inputs (see Figure 5). Due to severe overfitting and failure to generate usable outputs, the baseline could not achieve measurable Q1/Q2 alignment or SSIM scores. As a result, quantitative comparison metrics are not reported.
Figure 5: Failure cases from the instruction-tuned baseline. Although the test-time natural images (top row) are diverse and unseen during training, the model fails to adapt and instead memorizes training prompt-class pairs, producing incorrect tactile outputs (bottom row). Memorized outputs correspond to tomato, helicopter, and bed classes.
These findings highlight the challenge of applying instruction-tuned diffusion models directly for tactile graphic translation tasks, particularly when the model fails to balance prompt guidance with visual conditioning. Moreover, the relatively low pixel and color complexity of tactile images makes them poorly suited for standard full-model fine-tuning approaches, which typically require large, information-rich datasets to avoid overfitting. Our approach using LoRA and Dreambooth better mitigates these challenges by enabling more effective and data-efficient adaptation.
IV-B Fine-Tuning Configuration
@@
We fine-tuned Anything V3 [29] across 66 distinct classes. Each class had between 9 (llama) and 102 (helicopter) training images, averaging approximately 16 images per class. We leveraged sourced tactile images along with the prompts to develop class-specific adapters. These lightweight modules specialize in generating tactile graphics for each category and are integrated atop frozen versions of SD v1.5 for image generation. For fine-tuning, the Dreambooth configuration prioritized prior preservation with a loss weight of 1.0, ensuring the model retained its ability to generate generic images while adapting to tactile-specific features. The LoRA setup used a network module with linear dimensions set to 32 and an alpha parameter of 16, controlling the scaling of low-rank updates. This configuration enabled effective parameter adaptation without requiring pre-loaded weights.
IV-B1 Optimization and Hardware Configuration Parameters
@@
The optimization process employed the AdamW8bit optimizer [32], with a learning rate of $1\times 10^{-4}$ for the UNet [33] and $5\times 10^{-5}$ for the text encoder, using a constant rate scheduler without warm-up steps. Fine-tuning experiments were facilitated by NVIDIA Tesla T4 GPUs via Google Colab. Training used a batch size of 6 with mixed precision (FP16) across up to 20 training epochs.
IV-C Image Generation Configuration
@@
Text-to-image generation employed the DPM++ 2M Karras sampling method, using 20 steps at 512×512 pixel resolution and a CFG scale of 7. For image-to-image translation, configurations were similar but used a denoising strength of 0.9 to retain the structural essence of the original images while incorporating tactile features.
@@
For challenging classes, deviations were necessary to optimize structure and clarity. Objects requiring finer details—such as bee wings, elephant textures, and duck feathers—used higher denoising strengths (0.96–1.0) and adjusted CFG scales (8–10) for improved edge clarity. Classes with intricate internal features (e.g., basketball, dinosaur teeth, camera lenses) leveraged ControlNet lineart modules for contour refinement. Water-based objects (e.g., sailboat, floating duck) employed negative prompts to suppress extraneous water details. Simpler structures (e.g., egg, book, hatchback car) used LoRA-only adaptations without ControlNet assistance. Image generation was conducted on NVIDIA TITAN V GPUs (12288 MiB memory), with an average generation time of approximately 10 seconds per image.
IV-D Evaluation Protocol
@@
To assess the quality and usability of the generated tactile graphics, we developed a custom evaluation protocol in collaboration with tactile graphics designers. Conventional metrics such as Fréchet Inception Distance (FID) [34] and Structural Similarity Index (SSIM) [35] do not fully capture tactile-specific requirements like line clarity, texture, and adherence to accessibility guidelines [2]. Therefore, we prioritized expert evaluation by our industry partners, while also employing SSIM as a supplementary quantitative measure of structural fidelity.
IV-D1 Interface Design and Rationale
Our evaluation interface presented evaluators with pairs of images:
@@ IV-D1 Interface Design and Rationale
  • Tactile Image: Either a generated tactile graphic or a sourced tactile graphic from established libraries (e.g., APH, Perkins).
This side-by-side comparison allowed evaluators to assess how well the tactile graphic captured essential features and posture while adhering to tactile design principles. Evaluators were blind to the graphic’s source (generated vs. sourced), ensuring unbiased assessment. For evaluation, each class included one reference natural image and two tactile graphics (one generated, one sourced), totaling 66 natural images and 132 tactile graphics.
IV-D2 Evaluation Questions and Metrics
@@
Assess whether the tactile graphic follows established tactile graphics standards (e.g., BANA guidelines [2]).
@@
IV-D3 Natural Image Selection and Prompt Generation
@@
  • Original Prompt: The prompt generated during the prompt generation phase (Section III-B2).
  • Paraphrased Prompt: A refined version of the original prompt, paraphrased using DeepSeek [36] with the instruction: ”Paraphrase the given prompt for tactile generation while preserving the ’tactile’ subject and essential features.”
@@
IV-D4 Statistical Analysis Plan
@@
V RESULTS AND DISCUSSION
@@
TABLE I: Overview of the dataset statistics: scrapped tactile images (Source) used for training the models and the output tactile images (Generated) of our text-to-image adapters.
In this section, we present results including TactileNet dataset statistics (Table I) and outcomes from text-to-image and image-to-image translation tasks.
V-A Image-to-Image Translation Evaluation
Table II summarizes the quality ratings for both sourced and generated tactile images. Figure 2 illustrates sample outputs from our evaluation: the first row presents the natural reference images, the second row displays the sourced tactile graphics, and the third row showcases the tactile graphics generated using our fine-tuned adapters.
TABLE II: Quality Ratings for Generated vs. Sourced Tactile Graphics
Category Generated (%) Sourced (%)
Accept as Is 32.14 35.71
Accept with Minor Edits 39.29 39.29
Accept with Major Edits 28.57 21.43
Reject (Useless) 00.00 3.57
V-A1 Key Findings
Our evaluation reveals that both sourced and generated tactile graphics achieved 100% alignment with natural images in terms of pose and structural features (Q1), confirming their accuracy in representation. However, adherence to tactile accessibility guidelines (Q2) was slightly higher for sourced graphics (96.43%) compared to generated ones (92.86%). Additionally, 28.57% of generated graphics required major edits, primarily due to excessive visual complexity, such as unintended 3D perspectives in objects like Chair and Bed (Figure 2), which can hinder tactile interpretability. A small fraction of sourced tactile images (3.57%) were rejected, mostly due to human-induced mismatches in dataset curation—for example, incorrect natural image pairings, as seen in the camel species (Figure 6). This analysis highlights the strengths of AI-generated tactile graphics while emphasizing areas for improvement, particularly simplifying complex structures and refining dataset sourcing for higher semantic accuracy.
V-B Structural Fidelity and Abstraction Analysis
To quantify how well the generated graphics ($G$) preserve visual and tactile-relevant structure, we computed Structural Similarity Index (SSIM) scores across three comparisons: between generated images and tactile benchmarks ($G$ vs. $T$), tactile benchmarks and natural images ($T$ vs. $N$), and between both generated/tactile graphics and natural image binary masks ($G$ vs. $N_{\text{bin}}$, $T$ vs. $N_{\text{bin}}$).
TABLE III: SSIM Comparison Across Modalities
Comparison SSIM Interpretation
$G$ vs. $T$ 0.538 Model matches human tactile design fidelity
$T$ vs. $N$ 0.549 Human tactile abstraction baseline
$G$ vs. $N_{\text{bin}}$ 0.259 Model preserves silhouettes
$T$ vs. $N_{\text{bin}}$ 0.215 Human silhouette abstraction baseline
As shown in Table III, the similarity between $G$ and $T$ (0.538) nearly matches that of $T$ and $N$ (0.549), indicating that the model learns human-like abstraction patterns when generating tactile graphics. In silhouette comparisons using binary masks, $G$ outperformed $T$ (0.259 vs. 0.215), suggesting that the model retains object shape features more faithfully than traditional manual designs.
Figure 7: SSIM scores between generated ($G$), tactile benchmark ($T$), and natural ($N$) images across object classes. High-agreement classes (e.g., lamp, cup, banana) cluster in the upper right, while organic shapes (e.g., dog, bird) show greater divergence.
Class-wise analysis (Figure 7) highlights strong performance on structured objects like lamp, cup, and banana, while more organic or cluttered shapes (e.g., dog, bird) showed lower alignment. Overall, the SSIM analysis confirms that the model not only replicates human tactile design patterns but often improves upon them in preserving essential structural features.
V-C Text-to-Image Translation Results
@@
We generated a total of 32,000 tactile images across 66 classes, with 7,050 images retained after initial non-expert human filtering. Unlike the image-to-image translation setup, no natural image was provided as input; instead, the model relied solely on textual prompts to generate tactile graphics. Table I summarizes the per-class counts. Our fine-tuned adapters demonstrated adaptability through prompt-based edits (Figure 8), enabling modifications such as logo removal or pocket additions. The prompt used for generating images was: Base Prompt: Create a tactile graphic of a t-shirt with a pocket for the visually impaired, highlighting the round neckline, pocket edges, and hem with raised lines. Ensure the pocket is a distinct raised rectangle, allowing users to discover and feel the detail on the shirt’s chest.
Figure 8: Prompt editing for customization: (Left) T-shirt generated using the base prompt; (Center) Logo removed by adding “logo” as a negative prompt; (Right) Pocket removed by omitting the keyword “pocket” from the prompt.
VI Limitations and Future Work
VI-A Limitations
The current dataset is comprehensive but limited in size and diversity, with some classes having fewer images, which may hinder the model’s generalization, while the reliance on class-specific adapters restricts scalability, highlighting the need for a more efficient, generalized model for practical applications.
VI-B Future Work
Future efforts will focus on expanding TactileNet to include a diverse set of high-quality tactile images, enabling improved image-to-image translation systems that better manage complex tactile features. Incorporating generated images into the training pipeline will facilitate iterative model performance improvements, creating a self-reinforcing enhancement cycle. We aim to develop a single, generalized model capable of generating high-quality tactile graphics across all classes, leveraging our generated tactile images by employing techniques such as self-supervised training [37, 38]. Continued collaboration with tactile graphics designers and educators will ensure the generated graphics align with real-world usability standards and incorporate expert feedback into the training and evaluation process.
VII CONCLUSIONS
@@
This study marks a significant advancement in developing AI-driven accessibility tools through Generative AI, addressing the critical need for scalable, high-quality tactile graphics. Our key contributions are threefold:
@@
  • TactileNet Dataset: A first-of-its-kind collection of 1,029 expert-curated tactile images across 66 classes, enabling AI model training aligned with tactile design principles.
@@
  • Efficient Fine-Tuning: Integration of LoRA and DreamBooth reduced labor costs while achieving 92.86% adherence to tactile guidelines, rivalling manually sourced graphics.
@@
  • Human-Centric Evaluation: A protocol validated by tactile experts confirmed usability, with generated graphics requiring no rejections and achieving 100% alignment with natural images.
@@
By automating tactile graphic generation, our method accelerates production and democratizes access to educational materials for visually impaired learners. This work illustrates how AI can drive social good, offering a blueprint for inclusive technology that prioritizes human needs. Our framework demonstrates how scalable, AI-augmented tools can empower tactile designers and educators, promoting inclusive learning environments worldwide.
Acknowledgment
This work was supported in part by MITACS and the Digital Alliance of Canada. We also thank the student volunteers at the Intelligent Machines Lab (iML), Carleton University, for their valuable contributions.